OpenStudy (anonymous):

HELP with Geometry . 5 days behind .. Plans for a stadium are drawn on a coordinate grid. One wall lies on the line y = 2x +3 . A perpendicular wall passes through the point (4,-7). What is the equation of the line that passes through the new wall ? A . y = -1/2x -5 B. y = 2x-15 C. y = 1/2x -9 D . -2x +1

OpenStudy (anonymous):

A perpendicular slope is the opposite reciprocal to the existing slope of 2...the perpendicular line has a slope of -1/2 Using the slope intercept formula, y=mx+b, substitute in what you know (slope & point) -7=-1/2(4)+b -7 = -2 + b -5=b Using the perpendicular slope and b, write your equation: y=-1/2x-5

OpenStudy (anonymous):

Thank you @DSS I did my own work but this is as far as I got .. m1 x m2 = -1 2 x m2 = -2 m2 = -1/2 y + 7 = -1/2 (x-4) But I didn't know how to put it into standard form .

OpenStudy (anonymous):

ok I'll show you this way

OpenStudy (anonymous):

y + 7 = -1/2 (x-4) y + 7 = -1/2x + 2 y + 7-2 = -1/2x + 2-2 y+5 = -1/2x (-2)(y+5) = -1/2(-2)x -2y - 10 = x My answer says the same if you get rid of the fraction -2y=(-1/2x-5) -2y=x+10 -2y-10=x equivalent answers